OpenStudy (lena772):

Sam stores manure for his farm in a rectangular box. The length of the box is 7 feet, width is 9 inches, and height is 6 inches. Each bag of manure he bought contains 2.625 cubic feet of manure. How many bags of manure will fill the box completely? [1 foot = 12 inches]

OpenStudy (lena772):

i guess i would first find the area which would be 378 in^2 converting that to feet would give 31.5 sqft

OpenStudy (anonymous):

keep in mind length height & width are not in one unit you need to convert them first too

OpenStudy (lena772):

oh i totally didn't catch that part

OpenStudy (anonymous):

width = 0.75ft height = 0.5ft yes then you have to get the area

OpenStudy (lena772):

7*0.5=3sqft is the area volume is 2.625 cubed feet

OpenStudy (lena772):

so 1 bag of manure

OpenStudy (lena772):

how is volume ft ^2?

OpenStudy (anonymous):

ooops :P

OpenStudy (anonymous):

2.652 ft^3 my bad

OpenStudy (lena772):

so it's 1 bag of manure then?

OpenStudy (lena772):

this is a rectangular box

OpenStudy (lena772):

Surface Area = 2(lh) + 2(lw) + 2(wh) Volume = l x h x w

OpenStudy (anonymous):

oh lol okie sorry got confused :)

OpenStudy (anonymous):

yes so 1 bag of manure